How much is the power rating of a light bulb that uses 1.30x 107 J in 1 day?

a. 150 kW b. 150 hp c. 0.15 hp d. 0.15 k


1
Expert's answer
2021-04-23T07:27:07-0400

By definition, the power is:


"P = \dfrac{A}{t}"

where "A = 1.3\times 10^{7}J" and "t = 1day = 24\cdot 60\cdot60s = 86400s" is the time (number of seconds in one day).

Thus, obtain:


"P = \dfrac{1.3\times 10^{7}J}{86400s} \approx 150W = 0.15\space kW"

Answer. "0.15\space kW".